# MMformer with Adaptive Transferable Attention: Advancing Multivariate Time Series Forecasting for Environmental Applications

### v1

@@ 4.2 Evaluation Baseline Metrics
    Mean Squared Error (MSE) [69]: MSE measures the average of the squared differences between predicted and actual values, making it sensitive to larger errors. This sensitivity helps identify models with lower variance in their predictions.
@@
    Each error $(y_{i}-\hat{y}_{i})$ is squared in this formula, amplifying more significant errors. This means the MSE heavily penalizes significant deviations from the actual values, highlighting models with lower prediction variance.
@@
4.3 Application Analysis on Datasets
@@ 4.3.1 Basic Statistical Analysis of Data
Table 4 and Table 5 summarise the statistics of key indicators in the air quality and climate datasets.
@@
    The unit of Temperature: 0.1℃
Based on Table 4 analysis, all air quality features show minimum values of 0, with significant variance across all indicators except $CO$, indicating substantial regional air quality disparities. The consistently lower median values compared to means suggest right-skewed distributions, where generally good air quality is punctuated by severe pollution events. These spatial and temporal variations in pollution levels underscore the necessity for robust, targeted monitoring systems in high-risk areas.
From Table 5, 2013 precipitation data shows the highest IQR and variance, with a mean similar to 2012 but higher than 2014. 2014 had maximum precipitation, the most extensive range. Temperature means across years were comparable, though 2013 exhibited more significant variance and IQR, indicating more dispersed middle values.
4.3.2 Comparative Analysis of Prediction Methods in Air Quality Data
We compare the proposed MMformer with three baseline models—Transformer, iTransformer, and SARIMAX—using MSE, MAE, and MAPE as primary evaluation metrics. the Transformer is considered a benchmark model in MTS forecasting [71] because its self-attention mechanism. iTransformer refines the Transformer architecture to enhance robustness and accuracy in MTS and achieves state-of-the-art performance in real-world scenarios [72]. SARIMAX (Seasonal Autoregressive Integrated Moving Average Model with Exogenous Variables) is a widely adopted statistical model that offers theoretical interpretability by incorporating seasonal and exogenous factors [73]. Including SARIMAX highlights the contrast between neural networks and traditional methods in complex, multivariate problems.
Table 6: Multivariate time series forecasting results using a real-world air quality dataset. Results are averaged across all prediction lengths (Avg). Figures in bold are the best for that metric
@@ 4.3.2 Comparative Analysis of Prediction Methods in Air Quality Data
MMformer 0.53 0.48 26.53%
Transformer[57] 1.41 0.83 51.32%
iTransformer[21] 1.11 0.60 33.47%
SARIMAX[74] 215.17 6.90 10.86%
@@
Figure 11: Models Prediction Comparison in Air Quality Data
Table 6 demonstrates that, despite using a high-performance system (Intel Core i9-12900KF CPU, 64GB RAM, RTX 3090 GPU), the three neural network models substantially outperform SARIMAX across most metrics, except for MAPE. This suggests that traditional time series models struggle with variability in multivariate data. In contrast, deep networks may sometimes over- or underestimate extreme values, inflating relative errors and thus MAPE. However, while SARIMAX can avoid large percentage deviations when trends are more obvious, its global accuracy—reflected by MSE and MAE—remains lower than that of the neural network approaches.
For data preparation, we allocated 730 days for training, 366 days for validation, and 181 days for testing, employing a sliding window strategy to capture temporal dependencies. Sequence lengths, prediction horizons, and hyperparameters were carefully aligned across all models to ensure a fair comparison.
Figure 12: SARIMAX Prediction in Air Quality Test Dataset
Fig.11 compares all tested models except SARIMAX, which was excluded due to severe distortion that would hinder observation of other predictions. Although SARIMAX shows a respectable fit in Fig.12, it suffers from notable deviations at outlier points and abrupt changes, consistent with the inherent challenges that traditional methods face when modeling nonlinear structures. Fig.11 demonstrates MMformer’s overall advantage in capturing intrinsic properties of MTS data. Fig.16 – Fig.18 demonstrate each model’s predictions on the test dataset.
Figure 13: MMformer Prediction in Air Quality Test Dataset
Figure 14: iTransformer Prediction in Air Quality Test Dataset
Figure 15: Transformer Prediction in Air Quality Test Dataset
@@
We visualized the predictions for seven features on five test samples with the lowest MSE to further examine the models’ performance. MMformer, as shown in Fig.16, can more closely track the ground truth, even in significant variability. This suggests MMformer’s modeling structure and attention mechanisms more effectively capture global and local dependencies within the complex MTS.
@@
Figure 16: MMformer prediction in the test dataset with the five test samples with the lowest MSE in Air Quality
Figure 17: Transformer prediction in the test dataset with the five test samples with the lowest MSE in Air Quality
@@
Figure 18: iTransformer prediction in the test dataset with the five test samples with the lowest MSE in Air Quality
@@
In contrast, Fig.17 exhibits Transformer significant deviations from the truth, particularly in regions with abrupt fluctuations or sharp peaks. This suggests that the Transformer’s self-attention may struggle to maintain accurate predictions during periods of high variability in MTS. Fig.18 demonstrates the challenges faced by the iTransformer model when dealing with rapid changes in the data, which illustrates how predictions become distorted under significant variability. For example, in several subplots, sudden peaks in the truth are smoothed or delayed in the predictions. This suggests that embedding timestamps alone may be insufficient to maintain sequence integrity across diverse regions and may introduce higher sensitivity to noise or regional variability.
@@
In conclusion, our combined quantitative evaluation and systematic graphical analysis provide compelling evidence of the MMformer’s superior performance compared to the Transformer and iTransformer baselines. The observed shortcomings of the competing models in preserving critical patterns under adverse noise conditions further highlight the advantages of the proposed MMformer for modeling complex, multivariate time series.
4.3.3 Comparative Analysis of Prediction Methods in Climate Data
@@
Similar to the previous methodology, the climate dataset MMformer will be compared with Transformer, iTransformer, and traditional SARIMAX models using identical evaluation metrics. The analysis will assess actual prediction results to provide a comprehensive performance overview.
@@
Table 7: Multivariate time series forecasting results using a real-world climate dataset. Results are averaged across all prediction lengths (Avg). Figures in bold are the best for that metric
@@
MMformer 0.68 0.45 25.30%
Transformer 1.10 0.65 38.42%
iTransformer 0.74 0.45 25.64%
SARIMAX 12201.06 41.08 96.86%
Figure 19: Models Prediction Comparison in Climate Data
Figure 20: SARIMAX Prediction in Climate Test Dataset
Table 7 demonstrates that MMformer significantly outperforms other models when evaluated on comparable hardware, with all three neural network approaches substantially surpassing SARIMAX. This further confirms the limitations of traditional time series models. The dataset was partitioned into 160 days for training, 60 days for validation, and 56 days for testing, with sliding window techniques employed to maximize training data utilization. Sequence lengths and prediction horizons are the same across all models to ensure a fair comparison.
Figure 21: MMformer Prediction in Climate Test Dataset
Figure 22: iTransformer Prediction in Climate Test Dataset
Fig.20 demonstrates the limitations of the SARIMAX model, which significantly underperforms compared to the other methods across the entire time series. This poor performance on the dataset justifies the decision to exclude SARIMAX from the comparative analysis shown in Fig.19. The decline in SARIMAX performance when transitioning from air quality data to climate data can be attributed to fundamental characteristics of both the model and datasets. SARIMAX fundamentally depends on robustly identifying seasonal patterns and autoregressive relationships to generate accurate forecasts.
Figure 23: Transformer Prediction in Climate Test Dataset
Figure 24: MMformer prediction in the test dataset with the five test samples with the lowest MSE in Climate Dataset
Figure 25: iTransformer prediction in the test dataset with the five test samples with the lowest MSE in Climate Dataset
Figure 26: Transformer prediction in the test dataset with the five test samples with the lowest MSE in Climate Dataset
While the air quality dataset provides continuous observations over an extended timeframe (1277 days), the climate dataset contains only short temporal segments (276 days) of rainfall and temperature measurements collected exclusively during summer periods across three consecutive years. This temporal fragmentation disrupts the continuity necessary for effective seasonal pattern recognition. Furthermore, the substantially reduced time series length in the climate dataset compromises the statistical reliability of seasonal components and autoregressive relationships that SARIMAX relies upon for prediction accuracy.
The plots in Fig. 21 - 23 provide a detailed comparison of the model predictions against the actual target values for the test dataset. These figures illustrate the MMformer’s superior performance, which can closely track the fluctuations in the target variable over time. In contrast, Transformer and iTansformer exhibit more significant deviation from the actual values, particularly during periods of high volatility.
The consistent accuracy of MMformer’s predictions, as evidenced across these figures, supports the claim that it is the most effective MTS model for this application. The visual comparison of model outputs reinforces the quantitative metrics presented earlier, demonstrating MMformer’s superiority comprehensively over the other techniques evaluated.
@@
We visualized the predictions of the two features from the five test samples with the lowest MSE to further evaluate each model’s performance (Fig.2426). Fig.24 demonstrates that the MMformer model can accurately capture the overall trends and relationships between the features across the five test samples with the lowest MSE. The predicted values closely track the actual values, indicating that MMformer can preserve the global dependencies between the features even when the dataset dimensionality is reduced.
@@
In contrast, Fig.25 and Fig.26 reveal limitations in the ability of iTransformer and Transformer to model the local interactions between the two features jointly. While these models perform well on the reduced dataset according to the MSE metric in Table 7, the visualizations show that they struggle to capture the nuanced, localized patterns in the authentic feature relationships. The predicted values often deviate significantly from the ground truth, suggesting these models fail to preserve local dependencies when constrained to a fully constrained feature space.
By comparing the performance of the three models on these visualization plots, it is clear that the MMformer architecture is more effective at learning both the global trends and local interactions between features, even when the dataset dimensionality is greatly reduced. The quantitative metrics and visual analysis on the (331,1277,7) and (2415,276,2) datasets confirm that MMformer outperforms iTransformer, Transformer, and SARIMAX in modeling complex MTS relationships, demonstrating robust stability and generalization.

### v2

@@ 4.2 Evaluation Baseline Metrics
    Mean Squared Error (MSE) [60]: MSE measures the average of the squared differences between predicted and actual values, making it sensitive to larger errors. The sensitivity helps identify models with lower variance in their predictions.
@@
    Each error $(y_{i}-\hat{y}_{i})$ is squared in this formula, amplifying more significant errors. It means the MSE heavily penalizes significant deviations from the actual values, highlighting models with lower prediction variance.
@@
4.3 Results Analysis on Datasets
@@ 4.3.1 Basic Statistical Analysis of Data
Table 4 and Table 5 summarise the statistics of key indicators in the air quality and climate datasets. Based on Table 4 analysis, all air quality features show minimum values of 0, with significant variance across all indicators except $CO$, indicating substantial regional air quality disparities. The consistently lower median values compared to means suggest right-skewed distributions, where generally good air quality is punctuated by severe pollution events. These spatial and temporal variations in pollution levels underscore the necessity for robust, targeted monitoring systems in high-risk areas.
@@
Table 5 presents a statistical overview of precipitation and temperature data from 2018 to 2022. Precipitation means are comparable across these years, although annual variances and interquartile ranges (IQRs) show differences, with 2020 exhibiting the highest variance and 2018 the greatest IQR for precipitation. Temperature means are also consistent throughout the study period. However, temperature displayed more significant variability in 2018, as indicated by the highest variance and IQR, suggesting greater dispersion in middle temperature values. Overall, the analysis highlights a degree of interannual stability in these climate indicators.
@@
    The unit of Temperature: 1℃
4.3.2 Comparative Analysis of Prediction Methods on Real-World Data
Comparative Analysis on the Air Quality Dataset
We compare the proposed MMformer with four recent strong baseline models, iTransformer, PatchTST, TimesNet, and Transformer, using MSE, MAE, and MAPE as primary evaluation metrics. The reason for selecting these baseline models is that the ICLR 2024 Spotlight SOTA model, iTransformer, refines the Transformer architecture to enhance robustness and accuracy in MTS and achieves state-of-the-art performance in general MTS datasets [18]. PatchTST is recognized as a strong baseline for its innovative patching strategy, which divides time series into segments and applies a Transformer to capture both local dependencies within patches and global context across the series, demonstrating strong performance on diverse MTS benchmarks [21]. Similarly, TimesNet is highlighted for its unique approach to time series decomposition, effectively separating trend and seasonal components to better model complex temporal patterns, thereby achieving leading results on various benchmark tasks [22]. Because of its widely used self-attention mechanism, Transformer is considered one of the benchmark models in MTS forecasting [62]. Including these strong baseline models highlights the MMformer in environmental multi-regional MTS prediction.
Table 6: Performance Comparison of Multivariate Time Series Forecasting Models on the Air Quality Dataset. The seq_len and pred_len are set to 150 and 30 for all models. The results are rounded to three decimal places. Figures in red and bold are the best for that metric.
@@ 4.3.2 Comparative Analysis of Prediction Methods on Real-World Data
MMformer 0.245 0.332 18.872%
iTransformer 0.822 0.529 29.738%
TimesNet 0.807 0.528 29.319%
PatchTST 0.862 0.556 31.133%
Transformer 0.770 0.524 29.188%
Table 6 presents the results of MTS prediction on a real air quality dataset, evaluated using MSE, MAE, and MAPE metrics. Experimental results demonstrate that MMformer achieves the lowest MSE (0.245), MAE (0.332), and MAPE (18.872%), indicating a superior ability for error minimization and data fitting. A detailed comparative analysis reveals substantial improvements over all baseline methods. Compared to iTransformer, MMformer reduces MSE by 70.195%, MAE by 37.240%, and MAPE by 36.542%. Regarding TimesNet, MMformer decreases MSE by 69.641%, MAE by 37.121%, and MAPE by 35.632%. Compared to PatchTST, MMformer reduces MSE by 71.578%, MAE by 40.288%, and MAPE by 39.383%. Finally, relative to the Transformer model, the proposed MMformer shows significant reductions in MSE (decreased 68.182%), MAE (decreased 36.641%), and MAPE (decreased 35.343%). These findings strongly emphasize the MMformer’s robustness and effectiveness in predicting complex, high-dimensional air quality data, thus showcasing its substantial potential for practical environmental monitoring and forecasting applications.
The MMformer model incorporates MC Dropout for uncertainty quantification, thereby enabling the evaluation of prediction reliability. On the air quality dataset, MMformer exhibits a global average variance of 0.011919 and a global average standard deviation of 0.105612, reflecting the overall dispersion of its predictions. Across all samples, time steps, and features, the variance ranges from a minimum of 0.005076 to a maximum of 0.146630. The average 95% confidence interval width, estimated using the global average standard deviation, is approximately 0.4142 ($2\times 1.96\times 0.105612$). While the low global average variance and interval width suggest robust overall prediction confidence, the identified maximum variance of 0.146630 clearly indicates specific scenarios with reduced model confidence. Crucially for risk assessment, the maximum variance value directly flags the most unreliable prediction points as high-risk regions. In risk-sensitive applications, such as energy load forecasting or environmental disaster prediction, identifying these high-variance areas is paramount for effective risk management and optimized decision-making. Thus, MMformer’s variance analysis serves as a vital tool for reliability assessment and supports informed risk management strategies.
@@
Figure 6: MMformer, iTransformer, PatchTST, TimesNet, and Transformer Models Prediction Comparison on Air Quality Dataset
For data preparation, we allocated 730 days for training, 366 days for validation, and 181 days for testing, employing a sliding window strategy to capture temporal dependencies. Sequence lengths = 150, prediction lengths = 30, and hyperparameters are carefully aligned across all models to ensure a fair comparison. Fig.6 demonstrates MMformer’s overall advantage in capturing intrinsic properties of MTS data by the first feature sample of all baseline models. The visualization reveals that MMformer (orange line) most accurately tracks the true values (light blue line) throughout the prediction horizon, effectively capturing both gradual trends and sudden fluctuations. In contrast, other baseline models exhibit larger deviations from actual observations, particularly during periods of rapid change, highlighting MMformer’s superior capability in modeling complex climate patterns.
@@
We visualized the predictions for seven features on five test samples with the lowest MSE to further examine the models’ performance, as illustrated Figs.711. MMformer, as shown in Fig.7, the prediction effect is significantly better than that of baseline models. It suggests MMformer’s modeling structure and attention mechanisms more effectively capture global and local dependencies within the complex environmental MTS.
@@
Figure 7: MMformer prediction in the test dataset with the five test samples with the lowest MSE in the Air Quality Dataset
@@
Figure 8: iTransformer prediction in the test dataset with the five test samples with the lowest MSE in the Air Quality Dataset
Figure 9: PatchTST prediction in the test dataset with the five test samples with the lowest MSE in the Air Quality Dataset
Figure 10: TimesNet prediction in the test dataset with the five test samples with the lowest MSE in the Air Quality Dataset
Figure 11: Transformer prediction in the test dataset with the five test samples with the lowest MSE in the Air Quality Dataset
@@
In contrast, Fig.8 demonstrates the challenges faced by the iTransformer model when dealing with rapid changes in the data, which illustrates how predictions become distorted under significant variability. For example, in several subplots, sudden peaks in the truth are smoothed or delayed in the predictions. It suggests that embedding timestamps alone may be insufficient to maintain sequence integrity across diverse regions and may introduce higher sensitivity to noise or regional variability.
Figs.9 and 10 illustrate the prediction performance of PatchTST and TimesNet. These models effectively capture overall trends, with their predictions closely mirroring the actual data’s trajectory. However, they tend to exhibit smoothing and lagging when faced with rapid changes, spikes, or sharp turns in the time series, suggesting limitations in capturing high-frequency dynamics and sudden temporal shifts. Fig.11 exhibits Transformer significant deviations from the truth, particularly in regions with abrupt fluctuations or sharp peaks. It suggests that the Transformer’s self-attention may struggle to maintain accurate predictions during periods of high variability in MTS.
@@
In conclusion, based on the air quality dataset, our combined quantitative evaluation and systematic graphical analysis illustrate the MMformer’s superior performance. The observed shortcomings of the competing models in preserving critical patterns under adverse noise conditions further highlight the advantages of the proposed MMformer for modeling complex environmental MTS.
Comparative Analysis on the Climate Dataset
@@
In the climate dataset, MMformer will be compared with iTransformer, PatchTST, TimesNet, and Transformer using identical evaluation metrics. The analysis will assess actual prediction results and evaluation metrics to provide a comprehensive performance overview.
@@
Table 7: Performance Comparison of Multivariate Time Series Forecasting Models on the Climate Dataset. The seq_len and pred_len are set to 40 and 10 for all models. The results are rounded to three decimal places. Figures in red and bold are the best for that metric.
@@
MMformer 0.283 0.237 13.623%
iTransformer 0.578 0.342 18.878%
TimesNet 0.593 0.362 20.171%
PatchTST 0.634 0.388 21.304%
Transformer 0.411 0.399 23.776%
Table 7 demonstrates that MMformer also significantly outperforms other strong baseline models on the climate dataset, achieving the lowest Mean Squared Error (MSE) of 0.283, Mean Absolute Error (MAE) of 0.237, and Mean Absolute Percentage Error (MAPE) of 13.623%, which indicates excellent error minimization and data fitting capabilities. MMformer reduced MSE, MAE, and MAPE by 51.038%, 30.702%, and 27.837%, respectively, compared to iTransformer; by 52.277%, 34.530%, and 32.463% compared to TimesNet; and by 55.363%, 38.918%, and 36.054% compared to PatchTST. Furthermore, MMformer achieved even greater performance gains against the Transformer model, with reductions of 31.144% in MSE, 40.602% in MAE, and 42.703% in MAPE.
The MMformer model’s uncertainty quantification analysis on the climate dataset revealed a global average variance of 0.030902, corresponding to a global average standard deviation of 0.170853. It indicates the overall dispersion of the model’s predictions. The variance across all samples, timesteps, and features ranged from a minimum of 0.007938 to a maximum of 0.147739. Using a 95% confidence level (with a Z-score of approximately 1.96), the average width of the prediction confidence interval can be calculated as: $2\times 1.96\times 0.170853\approx 0.6690$. It allows for the evaluation of predictive confidence and the assessment of potential error margins in decision-making processes.
The dataset is partitioned into 1461 days for training, 181 days for validation, and the remaining days for testing, with sliding window techniques employed to maximize training data utilization. Fig.12 presents a prediction comparison of baseline models using a sample from the climate dataset. MMformer successfully captures these subsequent low actual values, thereby validating its adaptive capacity to varying temporal dynamics. The visual comparison further reinforces the quantitative metrics reported earlier, collectively demonstrating the comprehensive superiority of MMformer over the other evaluated models. The climate dataset, comprising five years (2018-2022, 1826 days) of continuous rainfall and temperature data, is selected for its lack of strong seasonal bias, which clearly demonstrates temporal variation robustness.
Figure 12: MMformer, iTransformer, PatchTST, TimesNet, and Transformer Models Prediction Comparison on the Cliamte Dataset
@@
To further assess model performance, we visualized the predictions for two features from the five test samples with the lowest MSE (Figs.13-17). As shown in Fig.13, MMformer accurately captures the overall trends and relationships between the selected features. The close alignment between predicted and actual values indicates that MMformer effectively maintains global dependencies between features, even when the dimensionality of the dataset is reduced.
Figure 13: MMformer prediction in the test dataset with the five test samples with the lowest MSE in the Climate Dataset
Figure 14: iTransformer prediction in the test dataset with the five test samples with the lowest MSE in the Climate Dataset
Figure 15: PatchTST prediction in the test dataset with the five test samples with the lowest MSE in the Climate Dataset
Figure 16: TimesNet prediction in the test dataset with the five test samples with the lowest MSE in the Climate Dataset
Figure 17: Transformer prediction in the test dataset with the five test samples with the lowest MSE in the Climate Dataset
@@
In contrast, Figs.14-17 reveal limitations in the ability of iTransformer, PatchTST, TimesNet, and Transformer to model the local interactions between the two features jointly. While these models perform well on the dataset according to the metrics in Table 7, the visualizations show that they struggle to capture the nuanced, localized patterns in the authentic feature relationships. The predicted values often deviate significantly from the ground truth, suggesting these models fail to capture complex environmental MTS local features.
By comparing the performance of these baseline models on metrics and visualization plots, it is clear that the MMformer is more effective at learning global trends and local interactions between features on real-world data. Whether it is an air quality dataset or a climate dataset, our combined quantitative evaluation and systematic graphical analysis provide compelling evidence of the MMformer’s superior performance compared to the iTransformer, PatchTST, TimesNet, and Transformer strong baseline models on real-world environmental datasets. To further evaluate MMformer’s generalization capabilities, we applied it to the widely adopted MTS datasets, PEMS.
Comparative Analysis of Prediction Methods in Widely Used MTS Datasets
To further validate MMformer’s generalization capabilities across diverse MTS benchmarks, we conducted a comparative analysis on the widely used PEMS datasets. The evaluation utilized four PEMS datasets (PEMS03, PEMS04, PEMS07, and PEMS08), with a sequence length of 40 and a prediction length of 10. MMformer is benchmarked against iTransformer, TimesNet, PatchTST, and Transformer, employing MSE, MAE, and MAPE as key evaluation metrics.
Table 8: Performance Comparison of Multivariate Time Series Forecasting Models on PEMS Datasets. The seq_len and pred_len are set to 40 and 10 for all models. The results are rounded to three decimal places. Figures in red and bold are the best for that metric, blue and bold indicate that metric is second.
Models MMformer iTransformer TimesNet PatchTST Transformer (Ours) (2024) (2023) (2023) (2017) Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE PEMS03 0.066 0.173 0.080 0.189 0.075 0.179 0.100 0.219 0.116 0.226 PEMS04 0.201 0.225 0.255 0.252 0.298 0.273 0.265 0.271 0.332 0.320 PEMS07 0.075 0.195 0.067 0.168 0.075 0.175 0.091 0.210 0.166 0.250 PEMS08 0.196 0.226 0.270 0.253 0.420 0.288 0.298 0.283 0.642 0.386
Table 9: Comparison of MAPE for Multivariate Time Series Forecasting Models on PEMS Datasets. The seq_len and pred_len are set to 40 and 10 for all models. The results are rounded to three decimal places. Figures in red and bold are the best for that metric, blue and bold indicate that metric is second.
Models MMformer iTransformer TimesNet PatchTST Transformer Metric MAPE MAPE MAPE MAPE MAPE PEMS03 9.801% 11.318% 10.706% 13.057% 12.997% PEMS04 12.803% 14.496% 15.584% 15.494% 17.706% PEMS07 11.428% 10.131% 10.524% 12.605% 14.055% PEMS08 12.684% 14.341% 16.138% 16.100% 19.908%
Table 10: Performance Comparison of Forecasting Baselines against the MMformer on PEMS Datasets. The seq_len and pred_len are set to 40 and 10 for all models. Results are rounded to three decimal places. Figures in blue and bold indicate MMformer is inferior to the model in this indicator.
Models Metric iTransformer (2024) TimesNet (2023) PatchTST (2023) Transformer (2017) MSE MAE MSE MAE MSE MAE MSE MAE PEMS03 +21.212% +9.249% +13.636% +3.468% +51.515% +26.590% +75.758% +30.636% PEMS04 +26.866% +12.000% +48.259% +21.333% +31.841% +20.444% +65.174% +42.222% PEMS07 -10.667% -13.846% 0.000% -10.256% +21.333% +7.692% +121.333% +28.205% PEMS08 +37.755% +11.947% +114.286% +27.434% +52.041% +25.221% +229.592% +70.796%
The experimental results, as detailed in Tables 8 through 10, indicate that MMformer consistently delivers superior performance across the evaluated PEMS datasets. The evaluation focuses on Mean Squared Error (MSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE).
On the PEMS03 dataset, MMformer exhibited the best performance, achieving the lowest MSE (0.066), MAE (0.173), and MAPE (9.801%). TimesNet followed, with an MSE of 0.075, MAE of 0.179, and MAPE of 10.706%. Compared to MMformer, iTransformer’s MSE and MAE increased by 21.212% and 9.249%, respectively. PatchTST shows increases of 51.515% in MSE and 26.590% in MAE, while Transformer’s MSE and MAE increased by 75.758% and 30.636%, respectively.
For the PEMS04 dataset, MMformer again demonstrated superior performance, obtaining the best MSE (0.201), MAE (0.225), and MAPE (12.803%). iTransformer ranked second with an MSE of 0.255, MAE of 0.252, and MAPE of 14.496%. Compared to MMformer, iTransformer’s MSE and MAE increased by 26.866% and 12.000%, respectively. PatchTST’s MSE and MAE increased by 31.841% and 20.444%, while Transformer’s MSE and MAE increased by 65.174% and 42.222%, respectively.
In the PEMS07 evaluation, iTransformer led in MSE (0.067), MAE (0.168), and MAPE (10.131%). MMformer also performed well, ranking third with an MSE of 0.075, MAE of 0.195, and MAPE of 11.428%. TimesNet secured the second position in MSE (0.075), MAE (0.175), and MAPE (10.524%). Compared to iTransformer, MMformer’s MSE and MAE increased by 10.667% and 13.846%, respectively. TimesNet’s MSE is comparable to MMformer’s, with an MAE increase of 10.256
MMformer’s performance is also strong on the PEMS08 dataset, achieving the best MSE (0.196), MAE (0.226), and MAPE (12.684%). iTransformer ranked second with an MSE of 0.270, MAE of 0.253, and MAPE of 14.341%. Compared to MMformer, iTransformer’s MSE and MAE increased by 37.755% and 11.947%, respectively. PatchTST’s MSE and MAE increased by 52.041% and 25.221%, while Transformer’s MSE and MAE increased by 229.592% and 70.796%, respectively.
Overall, MMformer demonstrated a consistently strong performance across the PEMS datasets. The experiments validate that MMformer is a viable and robust model for multivariate time series forecasting tasks, achieving SOTA or near-SOTA performance across multiple datasets and metrics. While iTransformer performs exceptionally well on the PEMS07 dataset, MMformer performs better on the other datasets. The highlights MMformer’s effectiveness and robustness in handling high-dimensional and complex MTS forecasting tasks, thus showcasing its substantial potential for environmental applications and other domains with similar data characteristics.